I am having trouble with my business math homework could someone please help me?
The question is Solve 5/8x + 3/5=8
I did ask one before but I forgot to put all the numbers in sorry.
I would like to thank you in advance for all the help you have giving me so far.

Answer by Theo(13342)

Your equation is:

5/(8x) + (3/5) = 8

Easiest thing to do is to get rid of all those denominators.

You can do this in the following way:

Multiply both sides of the equation by (5*8x) to get:

(5*8x) * (5/(8x)) + (5*8x) * (3/5) = (5*(8x)) * 8

Simplify to get:

25 + 24x = 320x

Substract 24x from both sides of this equation to get:

25 = 320x - 24x

Simplify to get:

25 = 296x

Divide both sides by 296 to get:

x = 25/296

Confirm your answer is good by substituting into the original equation to see if the equation is true.

Your original equation is:

5/(8x) + (3/5) = 8

Substitute 25/296 for x to get:

5/(8*25/296) + (3/5) = 8

Simplify to get:

5 / (200/296) + (3/5) = 8

This is equivalent to:

5 * (296/200) + (3/5) = 8

Simplify this to get:

(296/40) + 3/5) = 8

If you divide both numerator and denominator of (296/40) by 8, you will get:

(37/5) + (3/5) = 8 which reduces to:

8 = 8 confirming that x = 25/296 is good.

That's your answer.